A state's license plates consist of four letters followed by three numerals, and 240 letter arrangements are not allowed. How many plates can the state issue?
Answer by Edwin McCravy(20054)

A state's license plates consist of four letters followed by three numerals, and 240 letter arrangements are not allowed. How many plates can the state issue?
Choose the letters 26x26x26x26 - 240 = 456976 - 240 = 476736 ways

Choose the digits in 10x10x10 ways = 1000 ways,

Answer:  476736x1000 = 476736000

Edwin